OpenStudy (anonymous):

How do you solve sqr root 2x+4 - sqr root x =2?

myininaya (myininaya):

I would get one of the square root expressions by itself and square both sides.

myininaya (myininaya):

Then you would get the last square root expression by itself and square both sides one more time.

myininaya (myininaya):

And then you have no more square roots to worry with. Though you need to check your answers since you raised both sides to an even power.

OpenStudy (anonymous):

You have lost me :( this is my worst subject.... How do I get the square roots alone? Sorry

myininaya (myininaya):

By subtracting and/or adding to both sides.

myininaya (myininaya):

\[\sqrt{2x+4}-\sqrt{x}=2 \] Add sqrt(x) on both sides